# Stable Distributed Online Feedback Optimization for Distribution System Voltage Regulation

### v1

@@ IV Conclusion
In this letter, we focused on the distributed voltage regulation problem in distribution systems. We showed that the previously proposed two-metric approach might fail. As a remedy, we proposed a nested feedback optimization approach. Our simulation results showed that the approach achieved even better voltage regulation than a centralized approach while only requiring short-range communication between physical neighbours. The proposed distributed approach however does not extend to congestion management while the centralized does so. The approach works for either active or reactive power control, but not joint control. This is because the observation in Section II-D does not hold anymore since the voltage difference is no longer the gradient of either active or reactive power injection, but represents their joint non-separable impact. However, one might deploy active and reactive power feedback controllers separately at different sampling time. Whether the system under two feedback controllers is stable is left for future investigations.

### v2

@@ IV Conclusion
In this paper, we focused on the distributed voltage regulation problem in distribution systems and proposed a nested feedback optimization approach. We theoretically analyzed its convergence. Our simulation results showed that the approach, while only requiring short-range communication between physical neighbours, achieved satisfactory voltage regulation and outperformed other distributed approaches.
The proposed distributed approach does not extend to congestion management while the centralized does so. The current presentation is based on radial systems but the approach can be readily extended to meshed systems using the approach in [1] to construct the sensitivity matrix $\mathbf{X}$. Extending the approach to unbalanced systems is an interesting future direction, where the focus is to construct a positive definite $\mathbf{X}$ with a sparse inverse matrix. Finally, the approach works for either active or reactive power control, but not joint control. This is because the observation in Section II-D does not hold anymore since the voltage difference is no longer the gradient of either active or reactive power injection, but represents their joint non-separable impact. However, one might deploy active and reactive power feedback controllers separately at different sampling time. Whether the system under two feedback controllers is stable is left for future investigations.